OpenStudy (anonymous):

√36/√5 please help I give medals :) answer choices: A.6/5 B.6/√5 C.√5/5 D.6√5/5

OpenStudy (anonymous):

What is the square root of 36?

OpenStudy (tyteen4a03):

Hint: \(6^2 = 36\).

OpenStudy (anonymous):

6

OpenStudy (anonymous):

\[\sqrt{36}/\sqrt{5}\] 6sqrt(5) is your answer

OpenStudy (anonymous):

So, since it equals 6, you can substitute the 'square root of 36' with '6'

OpenStudy (anonymous):

Thanks for the help! :)

OpenStudy (anonymous):

the 6 came from 36 which is 6*6 and for every two of the same digits is one six, 5 is already simplified so it would be 6 sqrt(5)

OpenStudy (anonymous):

so b right?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

no problem :D glad I helped